Company has prepared department overhead budgets for budgeted-volume levels before allocations as follows:

Management has decided that the most appropriate inventory costs are achieved by using individual department overhead rates. These rates are developed after support-department costs are allocated to operating departments. Bases for allocation are to be selected from the following:

Support departments:

Building and grounds

$56,000

Personnel

1,140

General plant administration

14,040

Cafeteria: Operating loss

300

Storeroom

6,244

$77,724

Operating departments:

Machining

$34,000

Assembly

49,000

83,000

Total for support and operating departments

$160,724

Using the step-down method, allocate support-department costs. Allocate the costs of the support departments in the order given in this problem. (Round all indirect allocation rates to the nearest cent and all other computations to the nearest whole dollar. Use parentheses or a minus sign when decreasing departments by allocating costs.)

Support Departments

Operating Departments

General

Cafeteria

Bldg and

Plant

Operating

Grounds

Personnel

Admin.

Loss

Storeroom

Machining

Assembly

Costs incurred

Allocation of costs:

Building and grounds

Personnel

General plant administration

Cafeteria

Storeroom

Total budgeted costs of operating departments

Develop overhead rates per direct manufacturing labor-hour for machining and assembly using the step-down method. (Round the overhead rates to three decimal places.)

Step-down method:

Overhead rate per direct

Total costs allocated to department

/

Direct manufacturing labor-hours

=

manufacturing labor-hour

Machining

/

=

Assembly

/

=

Requirement 2. Using the direct method, rework requirement 1.

Using the direct method, allocate support-department costs. (Round cost allocations to the nearest whole dollar. Use parentheses or a minus sign when decreasing departments by allocating costs.)

Support Departments

Operating Departments

General

Cafeteria

Bldg and

Plant

Operating

Grounds

Personnel

Admin.

Loss

Storeroom

Machining

Assembly

Costs incurred

Allocation of costs:

Building and grounds

Personnel

General plant administration

Cafeteria

Storeroom

Total budgeted costs of operating departments

Develop overhead rates per direct manufacturing labor-hour for machining and assembly using the direct method. (Round the overhead rates to three decimal places.)

Direct method:

Overhead rate per direct

Total costs allocated to department

/

Direct manufacturing labor-hours

=

manufacturing labor-hour

Machining

$64,928

/

$

=

$

Assembly

$130,747

/

$

=

$

Requirement 3. Based on the following information about two jobs, determine the total overhead costs for each job by using rates developed in (a) requirement 1 and (b) requirement 2.

Start by computing the total overhead costs for each job using the rates developed in (a) requirement 1 (the step-down method.) (Round the total overhead costs to the nearest cent.)

(a) Step-down method:

Total

Overhead Costs

Job 88:

Machining

$

Assembly

$

Total cost of Job 88

$

Job 89:

Machining

$

Assembly

$

Total cost of Job 89

Now compute the total overhead costs for each job using the rates developed in (b) requirement 2 (the direct method). (Round the total overhead costs to the nearest cent.)

(b) Direct method:

Total

Overhead Costs

Job 88:

Machining

Assembly

Total cost of Job 88

Job 89:

Machining

Assembly

Total cost of Job 89


Requirements 1. Using the step-down method, allocate support-department costs. Develop overhead rates per direct manufacturing labor-hour for machining and assembly. Allocate the costs of the support departments in the order given in this problem. Use the allocation base for each support department you think is most appropriate 2. Using the direct method, rework requirement 1 3. Based on the following information about two jobs, determine the total overhead costs for each job by using rates developed in (a) requirement 1 and (b) requirement 2 4. The company evaluates the performance of the operating department managers on the basis of how well they managed their total costs, including allocated costs. As the manager of the Machining Department, which allocation method would you prefer from the results obtained in requirements 1 and 2? Explain PrintDone
Requirement 1. Using the step-down method, allocate support-department costs. Develop overhead rates per direct manufacturing labor-hour for machining and assembly. Allocate the costs of the support departments in the order given in this problem. Use the allocation base for each support department you think is most appropriate Begin by selecting the allocation base for each support department. Support Department Building and grounds Personne General plant administration Allocation Base Cafeteria: operating loss Storeroom
Answer #1
1 Step down method:
Support departments Production departments
Allocation basis Building and grounds Personnel General plant administration Cafeteria: Operating loss Storeroom Machining Assembly
Costs 56000 1140 14040 -300 6244 34000 49000
Building and grounds Square feet of floor space
Cost per square feet = 56000/400000 = 0.14
P: 9000*0.14; G:6000*0.14; C:6000*0.14; S:9600*0.14; M:48000*0.14; A:321400*0.14
-56000 1260 840 840 1344 6720 44996
Personnel Number of employees
Cost per employee = (1140+1260)/300 = 8
G: 8*60; C:30*8; S:12*8; M:120*8; A:78*8
-2400 480 240 96 960 624
General plant administration Manufacturing labour hours
Cost per manufacturing labour hour = 15360/19200 = 0.8
C:3000*0.8; S:2000*0.8; M:6000*0.8; A:8200*0.8
-15360 2400 1600 4800 6560
Cafeteria: Operating loss Square feet of flow space
Cost per square feet = 3180/(9600+48000+321400) = 0.008
S: 9600*0.08; M:48000*0.08; A:321400*0.08
-3180 81 403 2697
Storeroom Number of requisitions
Cost per requisition = 9365/5000 = 1.873
M:3000*1.873; A:2000*1.873
-9365 5619 3746
0 0 0 0 0 52502 107623
Overhead rate
Step-down method: Overhead rate per direct
Total costs allocated to department / Direct manufacturing labor-hours = manufacturing labor-hour
Machining 52502 / 4000 = 13.125
Assembly 107623 / 12000 = 8.969
2 Direct method
Support departments Production departments
Allocation basis Building and grounds Personnel General plant administration Cafeteria: Operating loss Storeroom Machining Assembly
Costs 56000 1140 14040 -300 6244 34000 49000
Building and grounds Square feet of floor space
Cost per square feet = 56000/(48000+321400) = 0.1515
M: 48000*0.1515; A:321400*0.1515
-56000 7277 48723
Personnel Number of employees
Cost per employee = 1140/(120+78) = 5.7575
M: 120*5.7575; A:78*5.7575
-1140 691 449
General plant administration Manufacturing labour hours
Cost per manufacturing labour hour = 14040/(6000+8200) = 0.9887
M: 6000*0.9887; A:8200*0.9887
-14040 5932 8108
Cafeteria: Operating loss Square feet of flow space
Cost per square feet = -300/(48000+321400) = 0.0008
M:48000*0.0008; A:321400*0.0008
300 -39 -261
Storeroom Number of requisitions
Cost per requisition = 6244/(3000+2000)=1.24
M: 3000*1.24; A:2000*1.24
-6244 3746 2498
0 0 0 0 0 51607 108517
Overhead rate
Direct method Overhead rate per direct
Total costs allocated to department / Direct manufacturing labor-hours = manufacturing labor-hour
Machining 51607 / 4000 = 12.902
Assembly 108517 / 12000 = 9.043
a) Step-down method: Total
Overhead Costs ($)
Job 88: Machining 52501.7
Assembly 107622.7
Total cost of Job 88 160124.5
b. Direct method
Job 88: Machining 51607
Assembly 108517
Total cost of Job 89 160124

As a manager, I would prefer step down method as the costs are well allocated to support departments and then to production departments